OpenStudy (beth12345):

how to find x when... sin30/x = sin90/720

OpenStudy (amistre64):

i assume thats the law of sines

OpenStudy (beth12345):

yup

OpenStudy (amistre64):

multiply both sides by x/720

OpenStudy (amistre64):

and then divide both sides by sin90

OpenStudy (amistre64):

or simply reciprocate and multiply thru by sin30

OpenStudy (beth12345):

so \[\sin 30 \times 720 = \sin 90 \times x\]

OpenStudy (anonymous):

I plugged it into my calculator and crossed multiplied and got -795.7. hope that helps(:

OpenStudy (amistre64):

so far yes

OpenStudy (amistre64):

divide off sin90 to get x all alone

OpenStudy (amistre64):

sin90=1 if we are talking in degrees

OpenStudy (beth12345):

is it then sin0.5x720=X ?

OpenStudy (amistre64):

nope; sin(30) x 720 ------------ = X sin(90)

OpenStudy (amistre64):

whats that sin0.5 about?

OpenStudy (beth12345):

i did sin30/sin90

OpenStudy (beth12345):

but with your equation, is the answer 360?

OpenStudy (amistre64):

looks that way as long as we are taking in terms of degrees and not radians

OpenStudy (beth12345):

ok, ya it would be degrees, thanks :)

OpenStudy (amistre64):

yw